# Stepsize anything: A unified learning rate schedule for budgeted-iteration training

### v1

@@ 1 Introduction
Deep learning has achieved remarkable success across various across a wide range of domains, including computer vision and natural language processing. However, despite continual advancements in hardware technologies [44, 52], the training cost of neural networks has increased dramatically due to the growing scale of models and datasets [4, 46, 8, 47]. As a result, resource constraints, including computational power, memory, energy consumption, and time budgets, are emerging as significant bottlenecks in the training process [40, 56]. These challenges highlight the pressing need for budgeted training, which aims to achieve optimal model performance under fixed hardware and limited time.
@@
To achieve this, we first bridge the gap by constructing a unified budget-aware training optimization framework, which incorporates the robustness to landscape curvature variations induced by data distribution, sampling, network architectures and optimization. Then we obtain numerical solutions by gradient projection methods. To eliminate the need for repeated numerical optimization when applying our method to different networks, we propose a universal parametric function that approximates numerical solutions. We nominate the resulting schedule Unified Budget-Aware (UBA) schedule. It requires tuning only a single hyper-parameter $\varphi$, reducing the overhead of per-network numerical optimization. Moreover, we establish a theoretical connection between $\varphi$ and the condition number, adding interpretation and optimization difficulty-aware theoretical grounding. Besides, we prove the convergence for different values of $\varphi$. These theoretical analysis along with empirical results offer practical guidelines for $\varphi$ selection. We evaluate UBA through comprehensive experiments across vision and language tasks, spanning diverse architectures and iteration budgets. Specifically, for vision tasks, the UBA schedule demonstrates consistent superiority over baselines across all evaluated datasets and model scales under varying training iterations. For language tasks, we validate the effectiveness of UBA through extensive benchmarks with OLMo model(36M, 73M, 150M and 300M parameters). Results show that UBA achieves state-of-the-art performance across on approximately half of the benchmarks, and consistently outperforms baselines on the average scores.
@@
    We perform experiments and demonstrate that UBA surpasses the commonly-used schedules across diverse vision and language tasks, spanning network architectures (e.g., ResNet, OLMo) and scales, under different training-iteration budgets.
@@ 2 Related work
Within the context of budgeted training, a key area that remains under-explored is achieving the best possible model performance within a fixed number of iterations, i.e. ‘budgeted-iteration training’. In this domain, learning rate scheduling based methods are particularly aligned with the objectives of budgeted-iteration training. Smith et al. [42] propose a new learning rate schedule, named cyclical learning rates (CLR). It improves accuracy in fewer iterations without tuning and is relevant to super-convergence phenomenon [43]. Li et al. [29] introduce an alternative setting of existing learning rate schedules for budgeted training. Chen et al. [5] propose the reflected exponential schedule (REX) via a profile and sampling fashion. Learning rate based approaches achieve robust and high-performing results under various lengths of training iterations, which corresponds to our purpose in this work, i.e. budgeted-iteration training [33]. Besides, this approach is plug-and-play, requiring no substantial alterations to the underlying model structure, making it readily adaptable to various deep network frameworks. Therefore, we explore the proper learning rate schedule to achieve budgeted-iteration training.
@@
The learning rate plays a pivotal role in controlling the optimization process during network training. The common scheme is the step decay schedule. A typical instance decreases the learning rate by a decaying scalar 0.1 after $50\%$ epochs and by a decaying scalar 0.01 after $75\%$ epochs [20]. Then Loshchilov et al. [32] observe that sharp decreases may prevent models from escaping local minima and propose the cosine schedule function, which is the most popular schedule for language model pretraining. Although some schedules include the CLR [42], REX [5], Warmup-Stable-Decay (WSD) [24] and schedule from multi-power law [33], there is no consensus on the optimal choice. In addition, the detail can be found in Appendix F, including some works focus on adaptive learning rate methods.
@@ 3.1 Finite optimization under limited training iterations
To design a one-size-fits-all learning rate schedule, we construct a robust optimization model of learning rates across varying training conditions (see more conceptual illustration in Appendix A). Specifically, we aim to guarantees minimal loss within constrained training iterations under the worst-case conditions, proposing a budget-iteration-aware framework for learning rate optimization.
Definition (Finite optimization): Let $f(W,D)$ denote the function parameterized by a given neural network with parameters $W\in\mathbb{R}^{N}$ on the dataset $D$, $\xi$ denotes the data sampling on the dataset $D$, and let $\mathcal{F}$ be a function class.Let $\mathcal{L}$ be the loss function. Let $\eta_{t}$ be the learning rate at the $t$-th iteration, $T$ be a maximum number of training iterations, and $t$ be the current learning step. A finite optimization is
@@
When setting $(K>2)$, UBA extends to a multi-phase formulation, which offers three potential advantages. (i) Hierarchical optimization: by partitioning the budget-aware optimization into multiple phases, each near a different local minima, it improves approximation accuracy and captures the dynamic features of the loss surface. (ii) It helps escape saddle points, (iii) It generalizes the single-phase method, allowing for future extensions. Notably, a single-phase approach $(K=2)$ remains effective, as the robust budget-aware model inherently derives the optimal schedule function, as shown in Figure 1. For consistency with common practice and to better isolate scheduling effects, this paper focuses primarily on the single-phase implementation. The multi-phase approach is also compared in Appendix, with its schedule shown in Figure 1.
We name the proposed schedule Unified Budget-Aware (UBA) schedule for reasons. The robust budget-aware model minimizes the loss function uniformly across the optimization trajectory, resulting in stable performance. UBA provides a reliable, unified choice for practitioners, eliminating the need for case-by-case baseline comparisons and delivering consistent superiority across datasets, architectures, and training budgets. Lastly, UBA can uniformly approximate the behavior of existing schedules through simple parameter adjustments.
@@ 3.2 Theoretical analysis
By Proposition 1, the hyper-parameter $\varphi$ is related to the optimization difficulty. In that case, the optimization difficulty is explicitly quantified as the condition number and $\varphi=2\kappa$, where $\kappa=\frac{\lambda_{u}^{(k)}}{\lambda_{l}^{(k)}}$. Furthermore, Proposition 2 shows that when $\varphi$ approaches 2, the proposed learning rate schedule converges to the standard cosine schedule. It means that, in regions where the optimization difficulty is not large (i.e., the curvature is relatively flat), setting $\varphi\approx 2$ allows the learning rate schedule to naturally reduce to the cosine form. Besides, our schedule can approximate the behavior of existing schedules (e.g.step decay, cosine annealing, cyclic schedule or Rex schedule) through simple parameter adjustments, shown in Table 1. The detail can be found in Appendix C. More importantly, it outperforms these schedules by training convergence and final accuracy, supporting results can be found in experiment sections 4.1,4.2 and Appendix E.5.
@@
Rex $\varphi=0.8$ OneCycle $\varphi=2,k\leftarrow k+1,T_{2}-T_{1}=$ pct__\__start step
@@
$\begin{split}&\mathbb{E}\left[\mathcal{L}(f(W_{t},\xi))-\mathcal{L}(f(\bar{W}^{(k)},\xi))\right]\\ \leq&\Bigl{(}\frac{4(T_{k+1}-T_{k})+(\varphi-2)\pi}{4(T_{k+1}-T_{k})+(\varphi-2)\pi(t-T_{k})}\Bigr{)}^{\tau}\cdot\exp\Bigl{(}-2\lambda_{i}^{(k)}\eta_{\min}(t-T_{k})\Bigr{)}\lambda_{u}^{(k)}\|W_{1+T_{k}}-\bar{W}^{(k)}\|_{2}^{2}\\ +&\sigma^{2}\sum\limits_{i=_{1+T_{k}}}^{t}\eta_{i}^{2}\sum\limits_{j=1}^{N}(\lambda_{j}^{(k)})^{2}\exp\Bigl{(}-2\lambda_{j}^{(k)}\eta_{\min}(t-i)\Bigr{)}\cdot\left(\frac{4(T_{k+1}-T_{k})+(\varphi-2)\pi(i-T_{k})}{4(T_{k+1}-T_{k})+(\varphi-2)\pi(t-T_{k})}\right)^{\tau}\\ \end{split}$ (7)
@@ 4 Experiment results
A key advantage of UBA lies in its critical parameter $\varphi$. While an optimal $\varphi$ can further improve model performance, we intentionally fix its value across all tasks and architectures to ensure fair evaluation. We fix $\varphi=5$ for SGD and $\varphi=0.5$ for AdamW (see learning rate variations in Figure 1). The rationale for these choices is systematically analyzed in our ablation study 4.3.
@@
Research on budgeted-iteration training remains limited, with most existing approaches focusing on learning rate scheduling strategies [43, 29, 5]. To provide a fair comparison, we adopt several widely used and empirically effective learning rate schedules as baselines: Step(SS) [15], Cosine(CS) [32], Cyclical (CLR) and OneCycle (1C) [42, 43], Budgeted training(BT) [29] and Reflected Exponential (REX) [5]. Details of these baselines are provided in Appendix E.1.
@@ 4.1 Experiments for Vision Classification Tasks
We evaluate our proposed UBA schedule on vision benchmarks (e.g., CIFAR10/100 and ImageNet) using different architectures (VGG16, ResNet18, ResNet34, ResNet50). In addition, we independently train models using fixed epoch budgets of 25%, 50%, and 100% of the maximum training epochs, without reusing or interpolating results from longer training runs. This setup ensures that each budget setting is evaluated in isolation, thereby preserving the integrity of comparisons across low and high training budgets. Table 2 presents the validation accuracy across different training budgets, comparing UBA against six baseline schedules (see detailed results in Appendix E.4).
@@
Schedule CIFAR10-ResNet18 CIFAR100-ResNet34 ImageNet–ResNet50
training budget (epoch(%)) training budget (epoch(%)) training budget (epoch(%))
75 150 300 75 150 300 75 150 300
(25%) (50%) (100%) (25%) (50%) (100%) (25%) (50%) (100%)
SS 92.41 93.90 93.94 70.85 75.58 77.28 74.84 76.96 78.10
CS 93.66 94.15 95.40 63.88 68.84 70.43 75.99 77.79 79.10
CLR 91.89 92.23 95.32 72.63 73.81 74.80 72.87 74.88 76.91
1C 94.36 95.02 95.48 73.72 75.53 77.90 75.28 77.37 78.79
BT 93.24 94.28 95.55 72.53 75.40 78.49 75.71 77.48 78.66
REX 94.79 94.96 95.59 73.12 75.48 77.99 75.28 77.03 78.46
UBA(ours) 94.54 95.26 95.74 74.57 76.68 78.97 76.00 77.99 79.32
@@ 4.2 Experiments for language models
From Table 3, UBA achieves state-of-the-art performance on approximately 50% of the benchmarks across all scales while the second-best schedule varies. Furthermore, it achieves consistent superior average performance among all baselines. It highlights the stable superiority and reliability of UBA, providing a default-strong choice. Moreover, it demonstrates significant improvements in SciQ-73M($+1.7$) and ARC-E-300M($+2.63$), highlighting its ability to enhance generalization across diverse benchmarks. Besides, in Figure 2, UBA consistently achieves lower training loss and validation loss throughout the training, indicating the efficient training ability and downstream performance enhancement. Notably, while task-specific tuning of $\varphi$ can further improve model performance, we intentionally fix its value across all tasks and architectures to ensure fair evaluation. Remarkably, even with this universal $\varphi$ setting, UBA consistently outperforms baselines on diverse benchmarks, demonstrating inherent robustness to varying benchmarks and model scales.
@@ 4.3 Ablation Study
We perform a sensitivity analysis of $\varphi\in\{0.25,0.5,1.0,2.5,5,10\}$ in equation (5) , which controls the variation speed of learning rate (see detailed setting, results and analyses in Appendix E.7). Experiments 11 show AdamW prefers smaller $\varphi$ while SGD requires larger $\varphi$. We attribute this phenomenon to the preconditioning effect of AdamW. As the relationship formalized in Proposition 1), smaller $\varphi$ is favored for low optimization difficult, while larger $\varphi$ is beneficial for difficult optimization. AdamW adapts the learning rate by scaling gradients with the second moment estimate $\sqrt{v_{t}}$, implicitly reducing the condition number of the optimization landscape. Thus the schedule with smaller $\varphi$ is preferred. In contrast, SGD lacks such adaptive mechanisms, thus a larger $\varphi$ is more suitable for this situation. This alignment between theory and experiment underscores the importance of tailoring $\varphi$ to the optimizer’s characteristics. Overall, $\varphi$ is related to a generalized optimization difficulty, where optimization precondition effect, datasets distribution and network architecture are all related to optimization difficulty.
@@ 5 Conclusion
In this paper, we construct a unified budget-aware training optimization framework that inherently adapts to landscape curvature variations, enhancing training robustness. From this optimization framework, we propose the Unified Budget-Aware (UBA) schedule. Extensive experiments demonstrate UBA consistently surpasses the commonly-used schedules across diverse vision and language tasks, spanning network architectures (e.g., ResNet, OLMo) and scales, under different training-iteration budgets.

### v4

@@ 1 Introduction
Deep learning has achieved remarkable success across a wide range of domains, including computer vision and natural language processing. However, despite continual advancements in hardware technologies [45, 53], the training cost of neural networks has increased dramatically due to the growing scale of models and datasets [4, 47, 8, 48]. As a result, resource constraints, including computational power, memory, energy consumption, and time budgets, are emerging as significant bottlenecks in the training process [41, 58]. These challenges highlight the pressing need for budgeted training, which aims to achieve optimal model performance under fixed hardware and limited time.
@@
To achieve this, we first bridge the gap by constructing a unified budget-aware training optimization framework, which incorporates the robustness to landscape curvature variations induced by data distribution, sampling, network architectures and optimization. The framework employs a min-max optimization approach to guarantee performance under worst-case conditions, since the minimization of learning rate is operated on an upper bound. Then we obtain numerical solutions by gradient projection methods. To eliminate the need for repeated numerical optimization when applying our method to different networks, we propose a universal parametric function that approximates numerical solutions. We nominate the resulting schedule Unified Budget-Aware (UBA) schedule. It requires tuning only a single hyper-parameter $\varphi$, reducing the overhead of per-network numerical optimization. Moreover, we establish a theoretical connection between $\varphi$ and the condition number, adding interpretation and optimization difficulty-aware theoretical grounding. Besides, we prove the convergence for different values of $\varphi$. These theoretical analysis along with empirical results offer practical guidelines for $\varphi$ selection. We evaluate UBA through comprehensive experiments across vision and language tasks, spanning diverse architectures and iteration budgets. Specifically, for vision tasks, the UBA schedule demonstrates consistent superiority over baselines across all evaluated datasets and model scales under varying training iterations. For language tasks, we validate the effectiveness of UBA through extensive benchmarks with OLMo model(36M, 73M, 150M and 300M parameters). Results show that UBA achieves state-of-the-art performance across on approximately half of the benchmarks, and consistently outperforms baselines on the average scores. Moreover, our performance improvements come with negligible computational overhead, which constitutes one of its key advantages.
@@
    We perform experiments and demonstrate that UBA surpasses the commonly-used schedules across diverse vision and language tasks, spanning network architectures (e.g., ResNet, OLMo) and scales, under different training-iteration budgets. Moreover, our performance improvements come with negligible computational overhead, which constitutes one of its key advantages.
@@ 2 Related work
Within the context of budgeted training, a key area that remains under-explored is achieving the best possible model performance within a fixed number of iterations, i.e. ‘budgeted-iteration training’. In this domain, learning rate scheduling based methods are particularly aligned with the objectives of budgeted-iteration training. Smith et al. [43] propose a new learning rate schedule, named cyclical learning rates (CLR). It improves accuracy in fewer iterations without tuning and is relevant to super-convergence phenomenon [44]. Li et al. [30] introduce an alternative setting of existing learning rate schedules for budgeted training. Chen et al. [5] propose the reflected exponential schedule (REX) via a profile and sampling fashion. Zhang et al. [57] introduce the finite horizon optimization framework and apply it to linear programming, which optimizes the algorithm performance under a strict iteration budget. Learning rate based approaches achieve robust and high-performing results under various lengths of training iterations, which corresponds to our purpose in this work, i.e. budgeted-iteration training [34]. Besides, this approach is plug-and-play, requiring no substantial alterations to the underlying model structure, making it readily adaptable to various deep network frameworks. Therefore, we explore the proper learning rate schedule to achieve budgeted-iteration training.
@@
The learning rate plays a pivotal role in controlling the optimization process during network training. The common scheme is the step decay schedule. A typical instance decreases the learning rate by a decaying scalar 0.1 after $50\%$ epochs and by a decaying scalar 0.01 after $75\%$ epochs [21]. Then Loshchilov et al. [33] observe that sharp decreases may prevent models from escaping local minima and propose the cosine schedule function, which is the most popular schedule for language model pretraining.
Pan et al. [36] propose Eigencurve, and demonstrate that minimax-optimal convergence can be achieved for quadratic objectives when the Hessian’s eigenvalue distribution exhibits substantial skewness. Defazio et al. [11] developed the theoretically-grounded framework for learning rate scheduling, establishing optimal linear decay schedules and refinements through rigorous mathematical derivation. Their work additionally provides the most extensive empirical evaluation of scheduling approaches to date. Although some schedules include the CLR [43], REX [5], Warmup-Stable-Decay (WSD) [25] and schedule from multi-power law [34], there is no consensus on the optimal choice. In addition, the detail can be found in Appendix F, including some works focus on adaptive learning rate methods.
@@ 3.1 Finite optimization under limited training iterations
To design a one-size-fits-all learning rate schedule, we construct a robust optimization model of learning rates across varying training conditions (see more conceptual illustration in Appendix A). Specifically, we aim to guarantee minimal loss within constrained training iterations under the worst-case conditions, proposing a budget-iteration-aware framework for learning rate optimization.
Definition (Finite optimization): Let $f(W,D)$ denote the function parameterized by a given neural network with parameters $W\in\mathbb{R}^{N}$ on the dataset $D$. Let $\xi$ denote the data sampling on the dataset $D$, and let $\mathcal{F}$ be a function class.Let $\mathcal{L}$ be the loss function. Let $\eta_{t}$ be the learning rate at the $t$-th iteration, $T$ be a maximum number of training iterations, and $t$ be the current learning step. A finite optimization is
@@
When setting $(K>2)$, UBA extends to a multi-phase formulation, which offers three potential advantages. (i) Hierarchical optimization: by partitioning the budget-aware optimization into multiple phases, each near a different local minima, it improves approximation accuracy and captures the dynamic features of the loss surface. (ii) It helps escape saddle points, (iii) It generalizes the single-phase method, allowing for future extensions. Notably, a single-phase approach $(K=2)$ remains effective, as the robust budget-aware model inherently derives the optimal schedule function, as shown in Figure 1. For consistency with common practice and to better isolate scheduling effects, this paper focuses primarily on the single-phase implementation. The multi-phase approach is also compared in Appendix E.9, with its schedule shown in Figure 1.
We name the proposed schedule Unified Budget-Aware (UBA) schedule for the following reasons. The robust budget-aware model minimizes the loss function uniformly across the optimization trajectory, resulting in stable performance. UBA provides a reliable, unified choice for practitioners, eliminating the need for case-by-case baseline comparisons and delivering consistent superiority across datasets, architectures, and training budgets. Lastly, UBA can uniformly approximate the behavior of existing schedules through simple parameter adjustments.
@@ 3.2 Theoretical analysis
By Proposition 1, the hyper-parameter $\varphi$ is related to the optimization difficulty. In that case, the optimization difficulty is explicitly quantified as the condition number and $\varphi=2\kappa$, where $\kappa=\frac{\lambda_{u}^{(k)}}{\lambda_{l}^{(k)}}$. Furthermore, Proposition 2 shows that when $\varphi$ approaches 2, the proposed learning rate schedule converges to the standard cosine schedule. It means that, in regions where the optimization difficulty is manageable (i.e., the curvature is relatively flat), setting $\varphi\approx 2$ allows the learning rate schedule to naturally reduce to the cosine form. Besides, our schedule can approximate the behavior of existing schedules (e.g.step decay, cosine annealing, cyclic schedule or Rex schedule) through simple parameter adjustments, shown in Table 1. The detail can be found in Appendix C. More importantly, it outperforms these schedules by training convergence and final accuracy, supporting results can be found in experiment sections 4.1,4.2 and Appendix E.5.
@@
Rex $\varphi=0.8$ OneCycle $\varphi=2,k\leftarrow k+1,T_{2}-T_{1}=$ pct_\_start step
@@
[…]
@@ 4 Experiment results
Our implementation strictly follows the PyTorch official API, i.e. it is inherited from the base class torch.optim.lr_scheduler.LRScheduler (Appendix E.3). It maintains full compatibility with all torch.optim optimizers. The complete production-ready code is publicly available (see Appendix E.3 for link).
A key advantage of UBA lies in its critical parameter $\varphi$. While an optimal $\varphi$ can further improve model performance, we intentionally fix its value across all tasks and architectures to ensure fair evaluation. We fix $\varphi$ for SGD and AdamW respectively (see learning rate variations in Figure 1). The rationale for these choices is systematically analyzed in our ablation study 4.3. Since each schedule has a different optimal learning rate in practice, to thoroughly address this concern, we conduct additional experiments to evaluate the performance of all schedules with different learning rates in Appendix E.8.
@@
Research on budgeted-iteration training remains limited, with most existing approaches focusing on learning rate scheduling strategies [44, 30, 5]. To provide a fair comparison, we adopt several widely used and empirically effective learning rate schedules as baselines: Step(SS) [16], Cosine(CS) [33], Cyclical (CLR) and OneCycle (1C) [43, 44], Budgeted training(BT, mathematically equivalent to linear decay) [30] and Reflected Exponential (REX) [5]. Details of these baselines are provided in Appendix E.1.
@@ 4.1 Experiments for vision classification tasks
We evaluate our proposed UBA schedule on vision benchmarks (e.g., CIFAR10/100 and ImageNet) using different architectures (ResNet18, ResNet34, ResNet50). In addition, we independently train models using fixed epoch budgets of 25%, 50%, and 100% of the maximum training epochs, without reusing or interpolating results from longer training runs. This setup ensures that each budget setting is evaluated in isolation, thereby preserving the integrity of comparisons across low and high training budgets. Table 2 presents the validation accuracy across different training budgets, comparing UBA against six baseline schedules (see detailed results in Appendix E.4).
@@
Dataset Network Method training budget (epoch(%))
75 (25%) 150(50%) 300(100%)
[…]
75 (25%) 150(50%) 300(100%)
[…]
75 (25%) 150(50%) 300(100%)
[…]
@@ 4.2 Experiments for language models
From Table 3, UBA achieves state-of-the-art performance on approximately 50% of the benchmarks across all scales while the second-best schedule varies. Furthermore, it achieves consistent superior average performance among all baselines. It highlights the stable superiority and reliability of UBA, providing a default-strong choice. Moreover, it demonstrates significant improvements in SciQ-73M(+1.7+1.7) and ARC-E-300M(+2.63+2.63), highlighting its ability to enhance generalization across diverse benchmarks. Besides, in Figure 2, UBA consistently achieves lower training loss and validation loss throughout the training, indicating the efficient training ability and downstream performance enhancement. Notably, while task-specific tuning of $\varphi$ can further improve model performance, we intentionally fix its value across all tasks and architectures to ensure fair evaluation. Remarkably, even with this universal $\varphi$ setting, UBA consistently outperforms baselines on diverse benchmarks, demonstrating inherent robustness to varying benchmarks and model scales.
@@ 4.3 Ablation study
We perform a sensitivity analysis of $\varphi\in\{0.25,0.5,1.0,2.5,5,10\}$ in equation (5) , which controls the variation speed of learning rate (see detailed setting, results and analyses in Appendix E.7).
Figure 11 show AdamW prefers smaller $\varphi$ while SGD requires larger $\varphi$. We attribute this phenomenon to the preconditioning effect of AdamW. As the relationship formalized in Proposition 1), smaller $\varphi$ is favored for low optimization difficult, while larger $\varphi$ is beneficial for difficult optimization. AdamW adapts the learning rate by scaling gradients with the second moment estimate $\sqrt{v_{t}}$, implicitly reducing the condition number of the optimization landscape. Thus the schedule with smaller $\varphi$ is preferred. In contrast, SGD lacks such adaptive mechanisms, thus a larger $\varphi$ is more suitable for this situation. This alignment between theory and experiment underscores the importance of tailoring $\varphi$ to the optimizer’s characteristics. Overall, $\varphi$ is related to a generalized optimization difficulty, where optimization precondition effect, datasets distribution and network architecture are all related to optimization difficulty.
@@ 5 Conclusion
In this paper, we construct a unified budget-aware training optimization framework that inherently adapts to landscape curvature variations, enhancing training robustness. From this optimization framework, we propose the Unified Budget-Aware (UBA) schedule. Extensive experiments demonstrate UBA consistently surpasses the commonly-used schedules across diverse vision and language tasks, spanning network architectures (e.g., ResNet, OLMo) and scales, under different training-iteration budgets. Moreover, our performance improvements come with negligible computational overhead, which constitutes one of its key advantages.